# TESSERACT: Eliminating Experimental Bias in Malware Classification across Space and Time (Extended Version)

### v1

@@ 3.1. Motivational Example
| Experimental setting | Training | Testing | Drebin (Arp et al., 2014) | DeepDrebin (Grosse et al., 2017) | Drebin (Arp et al., 2014) | DeepDrebin (Grosse et al., 2017) | ||||
@@
The data presented in Table 1 reveal that both Drebin and DeepDrebin exhibit significantly lower performance in realistic scenarios (highlighted in bold with a green background in the last row, corresponding to columns with 10% malware in testing) compared to settings similar to those described in prior works such as (Arp et al., 2014; Grosse et al., 2017) (highlighted in bold with a red background). This discrepancy can be attributed to unintentional experimental biases, which we elaborate on in the following sections.
3.2. Temporal Experimental Bias
@@
Concept drift is a problem that occurs in machine learning when a model becomes obsolete as the distribution of incoming data at test-time differs from that of training data, i.e., when the assumption does not hold that data is independent and identically distributed (i.i.d.) (Jordaney et al., 2017; Barbero et al., 2022). In the ML community, this problem is also known as dataset shift (Sugiyama et al., 2009). Time decay is the decrease in model performance over time caused by dataset shift.
@@
Concept drift in malware combined with similarities among malware within the same family causes k-fold cross validation (CV) to be positively biased, artificially inflating the performance of malware classifiers (Allix et al., 2015; Miller et al., 2016; Miller, 2015). K-fold CV is likely to include in the training set at least one sample of each malware family in the dataset, whereas new families will be unknown at training time in a real-world deployment. The all-black squares in Table 1 for 10-fold CV refer to each training/testing fold of the 10 iterations containing at least one sample from each time frame. The use of k-fold CV is widespread in malware classification research (Miller, 2015; Rahbarinia et al., 2017; Suarez-Tangil et al., 2017; Maggi et al., 2013; Dahl et al., 2013; Tahan et al., 2012; Markel and Bilzor, 2014; Laskov and Šrndić, 2011; Zhang et al., 2014; Curtsinger et al., 2011); while a useful mechanism to prevent overfitting (Bishop, 2006) or estimate the performance of a classifier in the absence of concept drift when the i.i.d. assumption holds (see considerations in § 4.4), it has been unclear how it affects the real-world performance of machine learning techniques with non-stationary data that are affected by time decay. Hence, in the first row of Table 1, we quantify the performance impact with 10-fold CV in the Android domain.
@@
The second row of Table 1 reports an experiment in which a classifier’s ability to detect both past and future objects is evaluated (Arp et al., 2014; Grosse et al., 2017). Although this characteristic is important, high performance should be expected from a classifier in such a scenario: if the classifier contains at least one variant of a past malware family, it will likely identify similar variants. We thus believe that experiments on the performance achieved on the detection of past malware can be misleading; therefore future research should focus on building malware classifiers that are robust against time decay.
@@
In the third row of Table 1, we identify a novel form of temporal bias that arises when goodware and malware correspond to different time periods, often due to originating from distinct data sources (e.g., as observed in (Mariconti et al., 2017)). The black and gray squares in Table 1 indicate that, even though the testing objects for malware come after the training objects, the time windows for goodware and malware do not overlap. This situation can lead to the classifier learning to distinguish between applications from different time periods rather than differentiating between goodware and malware, thereby resulting in artificially inaccurate performance. For instance, spurious features such as new API methods may exhibit strong discriminatory power simply because malicious applications predate the introduction of those APIs.
@@
Temporal inconsistency in data can result in Spurious Correlations. As described by (Arp et al., 2022), this refers to a scenario where a ML model forms shortcut patterns to differentiate classes based on artifacts that are irrelevant to the actual security issue. For instance, consider a malware detector that is trained on a dataset heavily dominated by certain malware families prevalent at the time. The model might inadvertently learn to identify features that are specific to these families rather than the broader characteristics of malware. Such features, although effective in the short-term, may become redundant as new malware families emerge, which were not present during the training phase. This leads to a temporal inconsistency in the model’s performance, as the model is tailored to the specific context of the training data and may not generalize well to future threats (Arp et al., 2022). Consequently, it is essential for designers to be mindful of and avoid potential temporal inconsistencies during the data collection and sampling process.
3.3. Spatial Experimental Bias
@@
4. Space-Time Aware Evaluation
@@ 4.1. Evaluation Constraints
We now formalize three constraints that must be enforced when dividing into and for a realistic setting that avoids spatio-temporal experimental bias (§ 3). While C1 was proposed in past work (Allix et al., 2015; Miller et al., 2016), we are the first to propose C2 and C3—which we show to be fundamental in § 4.4.
@@
where and . The same should hold for the training: although violating Equation 2 in the training data does not bias the evaluation, it may affect the sensitivity of the classifier to unrelated artifacts. Equation 2 has been violated in the past when goodware and malware have been collected from different time windows (e.g., (Mariconti et al., 2017))—if violated, the results are biased because the classifier may learn and test on artifactual behaviors that, for example, distinguish goodware from malware just by their different API versions.
@@
4.2. Time-aware Performance Metrics
@@
To facilitate the comparison of different time decay plots, we define a new metric, Area Under Time (AUT), the area under the performance curve over time. Formally, based on the trapezoidal rule (as in AUROC (Bishop, 2006)), AUT is defined as follows:
@@
4.3. Tuning Training Ratio
We propose a novel algorithm that allows for the adjustment of the training ratio when the dataset is imbalanced, in order to optimize a user-specified performance metric (-Score, Precision, or Recall) on the minority class, subject to a maximum tolerated error, while aiming to reduce time decay. The high-level intuition of the impact of changing is described in § 3.3. We also observe that ML literature has shown ROC curves to be misleading on highly imbalanced datasets (Davis and Goadrich, 2006; He and Garcia, 2009). Choosing different thresholds on ROC curves shifts the decision boundary, but (as seen in the motivating example of Figure 4) re-training with different ratios (as in our algorithm) also changes the shape of the decision boundary, better representing the minority class.
Our tuning algorithm is inspired by one proposed by Weiss and Provost (Weiss and Provost, 2003); they propose a progressive sampling of training objects to collect a dataset that improves AUROC performance of the minority class in an imbalanced dataset. However, they did not take temporal constraints into account (§ 3.2), and heuristically optimize only AUROC. Conversely, we enforce C1, C2, C3 (§ 4.1), and rely on AUT to achieve three possible targets for the malware class: higher -Score, higher Precision, or higher Recall. Also, we assume that the user already has a training dataset and wants to use as many objects from it as possible, while still achieving a good performance trade-off; for this purpose, we perform a progressive sub-sampling of the goodware class.
@@ 4.3. Tuning Training Ratio
The grid-search phase iterates over different values of , with a learning rate (e.g., ), and keeps as the value leading to the best performance, subject to the error constraint. To reduce the chance of discarding high-quality points while down-sampling goodware, we prioritize the most uncertain points (e.g., points close to the decision boundary of the SVM) (Settles, 2012). This approach differs from the one in the conference version of Tesseract (Pendlebury et al., 2019), where we relax the constraint on line 1 from () to a broader range of (). We find that imposing a stringent boundary is unnecessary when an error rate control is already in place (see algorithm 1, algorithm 1). Finally, the grid-search explores multiple values of and stores the best ones. To capture time-aware performance, we rely on AUT (§ 4.2), and the error rate is computed according to the target (see above). Tuning examples are in § 4.4.
4.4. Tesseract: Revealing Hidden Performance
Here, we show how our methodology can reveal hidden performance of Drebin (Arp et al., 2014) and DeepDrebin (Grosse et al., 2017), along with their robustness to time decay.
@@ 4.4. Tesseract: Revealing Hidden Performance
Figure 6 reports several performance metrics of the two selected classification approaches as point estimates over time. The -axis reports the testing slots in month granularity, whereas the -axis reports different scores between 0.0 and 1.0. The areas highlighted in blue correspond to the AUT() over the 48 testing months. The black dash-dotted horizontal lines represent the best -Score from the original works Arp et al.(Arp et al., 2014) and Papernot et al.(Grosse et al., 2017), corresponding to results obtained with 10 hold-out random splits for Drebin and DeepDrebin, respectively; both these settings are analogous to k-fold from a temporal bias perspective, and violate both C1 and C2. The red dashed horizontal lines correspond to 10-fold obtained on our dataset, which satisfies C3.
The figure illustrates that DeepDrebin’s performance aligns closely with that reported in the original study by Papernot et al.(Grosse et al., 2017), as indicated by the similarity between the red-dashed and black dash-dotted lines. In contrast, Drebin exhibits a slightly reduced performance compared to its original evaluation in Arp et al.(Arp et al., 2014). We believe the difference of performance is due to different data distribution. Both approaches show a similar trend of performance in the testing data. The 10-fold CV results are 0.85 and 0.84, with an AUT() of 0.50 and 0.49 for Drebin and DeepDrebin, respectively. Focusing on the malware side (highlighted lines), we observe that both -Scores and Recall rates initially decrease, followed by a modest recovery after Month 28. Precision remains stable in the first 24 months for both methods, then experiences a notable decline after Month 24 and generally remains lower thereafter. There is no obvious performance decay, apart from Month 39, from goodware’s side (dimmed lines).
@@
The table also highlights another important aspect: the need to investigate the causes behind the performance discrepancies at different granularities. Such an investigation could involve Explainable AI (XAI) techniques, which can provide insights into the decision-making processes of classifiers (Chow et al., 2023). Understanding why a classifier’s performance varies with time can inform improvements in classifier design and contribute to developing more robust malware detection systems.
@@
Robustness of deep learning approach. In our initial analysis documented in the original paper (refer to the ‘DL’ plot of Figure 5 in (Pendlebury et al., 2019)), the deep learning method appeared to outshine the other two traditional ML approaches (Alg1(Drebin) and Alg2(MaMaDroid)) in terms of robustness after enforcing C1, C2 and C3. However, upon closer inspection and a more extended evaluation, this claim does not hold consistently. It is true that during the first two year of evaluation, the deep learning approach demonstrated a better and relatively stable performance. Nonetheless, when considering the broader scope of the analysis, it becomes evident that the overall performance of the deep learning method aligns closely with that of traditional machine learning techniques, if retraining is not possible. This revelation underscores that while deep learning may offer advantages in certain temporal slices, its efficacy is comparable to traditional models when viewed across a comprehensive timeline. This unresolved inquiry is delegated to future investigative efforts to discern the conditions under which DeepDrebin, or deep learning in general, exhibits superior performance within certain testing intervals.
@@
Tuning algorithm. We now evaluate whether our tuning algorithm (Algorithm 1 in § 4.3) improves robustness to time decay of a malware classifier for a given target performance. We first aim to maximize -Score of malware class, subject to . After running Algorithm 1 on Drebin (Arp et al., 2014) and DeepDrebin (Grosse et al., 2017), we set for Drebin, and for DeepDrebin. Figures 6(a) and 6(d) report the respective improvements in test performance of the two approaches on four years of testing data after applying to the full training set .
We remark that the choice of uses only training information (see Algorithm 1) and no test information is used—the optimal value is chosen from a 4-month validation set extracted from the one year of training data; this is to simulate a realistic deployment setting in which we have no a priori information about testing. 6(a) (Drebin) and 6(d) (DeepDrebin) show that our approach for finding the best improves the -Score on malware at most of the four-year testing time, at the cost of slightly reduced goodware performance. Table 3 shows details of how total FPs, total FNs, and AUT changed by training Drebin and DeepDrebin with , , and instead of . These training ratios have been computed subject to for , for , and for ; the difference in the maximum tolerated errors is motivated by the class imbalance in the dataset—which causes lower FPR and higher FNR values (see definitions in § 4.3), as there are many more goodware than malware.
@@
Furthermore, a three-year evaluation window was also employed to assess the tuning algorithm, as evidenced in Figures 6(b), 6(c) for Drebin, and Figures 6(e), 6(f) for DeepDrebin. This shorter period of evaluation suggests the necessity for regular updates to the tuning process to keep pace with the dynamic nature of malware threats. Aligning with the full dataset protocol, the first year of data from this three-year period was used to ascertain the optimal , and subsequent tests were conducted over the following two years (as listed in each sub-captions). The optimal was found to be 0.35 on both years 2015 and 2016 for both Drebin and DeepDrebin. Compared to the baseline performance (indicated by gray lines in the graphs), the tuned outcomes demonstrate enhanced robustness in the latter two years, comparing to in the long-run evaluation.
These results highlight the effectiveness of employing a short-term evaluation interval for both tuning and testing our algorithm. The graphical analyses, as demonstrated in Figures 6(b), 6(c), 6(e), and 6(f), reveal that our tuning algorithm not only enhances the overall robustness of the malware classifiers (Drebin and DeepDrebin) but also contributes to more consistent performance over time. This is particularly evident in the post-tuning improvements in the -Score, Precision, and Recall metrics across different time frames. The tuning process, guided by Algorithm 1, effectively addresses the challenge of time decay, as shown by the improved AUT metrics over various periods.

### v2

@@ 3.1 Motivational Example
| Experimental setting | Training | Testing | Drebin [6] | DeepDrebin [33] | Drebin [6] | DeepDrebin [33] | ||||
@@
The data presented in Table 1 reveal that both Drebin and DeepDrebin exhibit significantly lower performance in realistic scenarios (highlighted in bold with a green background in the last row, corresponding to columns with 10% malware in testing) compared to settings similar to those described in prior works such as [6, 33] (highlighted in bold with a red background). This discrepancy can be attributed to unintentional experimental biases, which we elaborate on in the following sections.
3.2 Temporal Experimental Bias
@@
Concept drift is a problem that occurs in machine learning when a model becomes obsolete as the distribution of incoming data at test-time differs from that of training data, i.e., when the assumption does not hold that data is independent and identically distributed (i.i.d.) [37, 8]. In the ML community, this problem is also known as dataset shift [66]. Time decay is the decrease in model performance over time caused by dataset shift.
@@
Concept drift in malware combined with similarities among malware within the same family causes k-fold cross validation (CV) to be positively biased, artificially inflating the performance of malware classifiers [2, 48, 49]. K-fold CV is likely to include in the training set at least one sample of each malware family in the dataset, whereas new families will be unknown at training time in a real-world deployment. The all-black squares in Table 1 for 10-fold CV refer to each training/testing fold of the 10 iterations containing at least one sample from each time frame. The use of k-fold CV is widespread in malware classification research [49, 55, 65, 44, 18, 67, 47, 38, 74, 17]; while a useful mechanism to prevent overfitting [11] or estimate the performance of a classifier in the absence of concept drift when the i.i.d. assumption holds (see considerations in § 4.4), it has been unclear how it affects the real-world performance of machine learning techniques with non-stationary data that are affected by time decay. Hence, in the first row of Table 1, we quantify the performance impact with 10-fold CV in the Android domain.
@@
The second row of Table 1 reports an experiment in which a classifier’s ability to detect both past and future objects is evaluated [6, 33]. Although this characteristic is important, high performance should be expected from a classifier in such a scenario: if the classifier contains at least one variant of a past malware family, it will likely identify similar variants. We thus believe that experiments on the performance achieved on the detection of past malware can be misleading; therefore future research should focus on building malware classifiers that are robust against time decay.
@@
In the third row of Table 1, we identify a novel form of temporal bias that arises when goodware and malware correspond to different time periods, often due to originating from distinct data sources (e.g., as observed in [46]). The black and gray squares in Table 1 indicate that, even though the testing objects for malware come after the training objects, the time windows for goodware and malware do not overlap. This situation can lead to the classifier learning to distinguish between applications from different time periods rather than differentiating between goodware and malware, thereby resulting in artificially inaccurate performance. For instance, spurious features such as new API methods may exhibit strong discriminatory power simply because malicious applications predate the introduction of those APIs.
@@
Temporal inconsistency in data can result in Spurious Correlations. As described by [5], this refers to a scenario where a ML model forms shortcut patterns to differentiate classes based on artifacts that are irrelevant to the actual security issue. For instance, consider a malware detector that is trained on a dataset heavily dominated by certain malware families prevalent at the time. The model might inadvertently learn to identify features that are specific to these families rather than the broader characteristics of malware. Such features, although effective in the short-term, may become redundant as new malware families emerge, which were not present during the training phase. This leads to a temporal inconsistency in the model’s performance, as the model is tailored to the specific context of the training data and may not generalize well to future threats [5]. Consequently, it is essential for designers to be mindful of and avoid potential temporal inconsistencies during the data collection and sampling process.
3.3 Spatial Experimental Bias
@@
4 Space-Time Aware Evaluation
@@ 4.1 Evaluation Constraints
We now formalize three constraints that must be enforced when dividing into and for a realistic setting that avoids spatio-temporal experimental bias (§ 3). While C1 was proposed in past work [2, 48], we are the first to propose C2 and C3—which we show to be fundamental in § 4.4.
@@
where and . The same should hold for the training: although violating Equation 2 in the training data does not bias the evaluation, it may affect the sensitivity of the classifier to unrelated artifacts. Equation 2 has been violated in the past when goodware and malware have been collected from different time windows (e.g., [46])—if violated, the results are biased because the classifier may learn and test on artifactual behaviors that, for example, distinguish goodware from malware just by their different API versions.
@@
4.2 Time-aware Performance Metrics
@@
To facilitate the comparison of different time decay plots, we define a new metric, Area Under Time (AUT), the area under the performance curve over time. Formally, based on the trapezoidal rule (as in AUROC [11]), AUT is defined as follows:
@@
4.3 Tuning Training Ratio
We propose a novel algorithm that allows for the adjustment of the training ratio when the dataset is imbalanced, in order to optimize a user-specified performance metric (-Score, Precision, or Recall) on the minority class, subject to a maximum tolerated error, while aiming to reduce time decay. The high-level intuition of the impact of changing is described in § 3.3. We also observe that ML literature has shown ROC curves to be misleading on highly imbalanced datasets [21, 36]. Choosing different thresholds on ROC curves shifts the decision boundary, but (as seen in the motivating example of Figure 4) re-training with different ratios (as in our algorithm) also changes the shape of the decision boundary, better representing the minority class.
Our tuning algorithm is inspired by one proposed by Weiss and Provost [71]; they propose a progressive sampling of training objects to collect a dataset that improves AUROC performance of the minority class in an imbalanced dataset. However, they did not take temporal constraints into account (§ 3.2), and heuristically optimize only AUROC. Conversely, we enforce C1, C2, C3 (§ 4.1), and rely on AUT to achieve three possible targets for the malware class: higher -Score, higher Precision, or higher Recall. Also, we assume that the user already has a training dataset and wants to use as many objects from it as possible, while still achieving a good performance trade-off; for this purpose, we perform a progressive sub-sampling of the goodware class.
@@ 4.3 Tuning Training Ratio
The grid-search phase iterates over different values of , with a learning rate (e.g., ), and keeps as the value leading to the best performance, subject to the error constraint. To reduce the chance of discarding high-quality points while down-sampling goodware, we prioritize the most uncertain points (e.g., points close to the decision boundary of the SVM) [60]. This approach differs from the one in the conference version of Tesseract [54], where we relax the constraint on line 1 from () to a broader range of (). We find that imposing a stringent boundary is unnecessary when an error rate control is already in place (see algorithm 1, algorithm 1). Finally, the grid-search explores multiple values of and stores the best ones. To capture time-aware performance, we rely on AUT (§ 4.2), and the error rate is computed according to the target (see above). Tuning examples are in § 4.4.
4.4 Tesseract: Revealing Hidden Performance
Here, we show how our methodology can reveal hidden performance of Drebin [6] and DeepDrebin [33], along with their robustness to time decay.
@@ 4.4 Tesseract: Revealing Hidden Performance
Figure 6 reports several performance metrics of the two selected classification approaches as point estimates over time. The -axis reports the testing slots in month granularity, whereas the -axis reports different scores between 0.0 and 1.0. The areas highlighted in blue correspond to the AUT() over the 48 testing months. The black dash-dotted horizontal lines represent the best -Score from the original works Arp et al.[6] and Papernot et al.[33], corresponding to results obtained with 10 hold-out random splits for Drebin and DeepDrebin, respectively; both these settings are analogous to k-fold from a temporal bias perspective, and violate both C1 and C2. The red dashed horizontal lines correspond to 10-fold obtained on our dataset, which satisfies C3.
The figure illustrates that DeepDrebin’s performance aligns closely with that reported in the original study by Papernot et al.[33], as indicated by the similarity between the red-dashed and black dash-dotted lines. In contrast, Drebin exhibits a slightly reduced performance compared to its original evaluation in Arp et al.[6]. We believe the difference of performance is due to different data distribution. Both approaches show a similar trend of performance in the testing data. The 10-fold CV results are 0.85 and 0.84, with an AUT() of 0.50 and 0.49 for Drebin and DeepDrebin, respectively. Focusing on the malware side (highlighted lines), we observe that both -Scores and Recall rates initially decrease, followed by a modest recovery after Month 28. Precision remains stable in the first 24 months for both methods, then experiences a notable decline after Month 24 and generally remains lower thereafter. There is no obvious performance decay, apart from Month 39, from goodware’s side (dimmed lines).
@@
The table also highlights another important aspect: the need to investigate the causes behind the performance discrepancies at different granularities. Such an investigation could involve Explainable AI (XAI) techniques, which can provide insights into the decision-making processes of classifiers [16]. Understanding why a classifier’s performance varies with time can inform improvements in classifier design and contribute to developing more robust malware detection systems.
@@
Robustness of deep learning approach. In our initial analysis documented in the original paper (refer to the ‘DL’ plot of Figure 5 in [54]), the deep learning method appeared to outshine the other two traditional ML approaches (Alg1(Drebin) and Alg2(MaMaDroid)) in terms of robustness after enforcing C1, C2 and C3. However, upon closer inspection and a more extended evaluation, this claim does not hold consistently. It is true that during the first two year of evaluation, the deep learning approach demonstrated a better and relatively stable performance. Nonetheless, when considering the broader scope of the analysis, it becomes evident that the overall performance of the deep learning method aligns closely with that of traditional machine learning techniques, if retraining is not possible. This revelation underscores that while deep learning may offer advantages in certain temporal slices, its efficacy is comparable to traditional models when viewed across a comprehensive timeline. This unresolved inquiry is delegated to future investigative efforts to discern the conditions under which DeepDrebin, or deep learning in general, exhibits superior performance within certain testing intervals.
@@
Tuning algorithm. We now evaluate whether our tuning algorithm (Algorithm 1 in § 4.3) improves robustness to time decay of a malware classifier for a given target performance. We first aim to maximize -Score of malware class, subject to . After running Algorithm 1 on Drebin [6] and DeepDrebin [33], we set for Drebin, and for DeepDrebin. Figures 7(a) and 7(d) report the respective improvements in test performance of the two approaches on four years of testing data after applying to the full training set .
We remark that the choice of uses only training information (see Algorithm 1) and no test information is used—the optimal value is chosen from a 4-month validation set extracted from the one year of training data; this is to simulate a realistic deployment setting in which we have no a priori information about testing. 7(a) (Drebin) and 7(d) (DeepDrebin) show that our approach for finding the best improves the -Score on malware at most of the four-year testing time, at the cost of slightly reduced goodware performance. Table 3 shows details of how total FPs, total FNs, and AUT changed by training Drebin and DeepDrebin with , , and instead of . These training ratios have been computed subject to for , for , and for ; the difference in the maximum tolerated errors is motivated by the class imbalance in the dataset—which causes lower FPR and higher FNR values (see definitions in § 4.3), as there are many more goodware than malware.
@@
Furthermore, a three-year evaluation window was also employed to assess the tuning algorithm, as evidenced in Figures 7(b), 7(c) for Drebin, and Figures 7(e), 7(f) for DeepDrebin. This shorter period of evaluation suggests the necessity for regular updates to the tuning process to keep pace with the dynamic nature of malware threats. Aligning with the full dataset protocol, the first year of data from this three-year period was used to ascertain the optimal , and subsequent tests were conducted over the following two years (as listed in each sub-captions). The optimal was found to be 0.35 on both years 2015 and 2016 for both Drebin and DeepDrebin. Compared to the baseline performance (indicated by gray lines in the graphs), the tuned outcomes demonstrate enhanced robustness in the latter two years, comparing to in the long-run evaluation.
These results highlight the effectiveness of employing a short-term evaluation interval for both tuning and testing our algorithm. The graphical analyses, as demonstrated in Figures 7(b), 7(c), 7(e), and 7(f), reveal that our tuning algorithm not only enhances the overall robustness of the malware classifiers (Drebin and DeepDrebin) but also contributes to more consistent performance over time. This is particularly evident in the post-tuning improvements in the -Score, Precision, and Recall metrics across different time frames. The tuning process, guided by Algorithm 1, effectively addresses the challenge of time decay, as shown by the improved AUT metrics over various periods.